OpenStudy (anonymous):

log base 3 X^2 - log base 3 x = 2

OpenStudy (anonymous):

\[\log_{3}X^2 - \log_{3}x = 2 \]

OpenStudy (akshay_budhkar):

lets solve it together..

OpenStudy (akshay_budhkar):

do you know log(a)^b=b*log(a)?

OpenStudy (akshay_budhkar):

you there?

OpenStudy (akshay_budhkar):

iseemc0nfused?

OpenStudy (akshay_budhkar):

you dun seem to want to solve it.. lol neways.. refer this http://www.wolframalpha.com/input/?i=log+base+3+X^2+-+log+base+3+x+%3D+2 all the best:)

OpenStudy (anonymous):

sorry i was away from the comp. but thanks :)

OpenStudy (anonymous):

i forgot the quotient rule!! =X taking a 4 year break from math was a bad idea

OpenStudy (anonymous):

i forgot the quotient rule!! =X taking a 4 year break from math was a bad idea

OpenStudy (anonymous):

here log base 3 = log given logx^2 -logx =2 logx^2=> 2logx 2logx - logx =2 logx=2 x = 3^2 = 9

OpenStudy (anonymous):

\[\log_{3}x^2=2\log_{3}x \] hence,\[\log_{3}x=2 and [\log_{3}x=( \log_{10}x)/\log_{10}3 \]

OpenStudy (anonymous):

x = 3^2 = 9